# Observing structural disorder induced interacting topological phase in an atom array

### v1

@@ Abstract
Topological phases of matter can appear in noninteracting systems 1, 2, 3, as in band topology, or interacting systems 4, such as in spin models, with their defining features typically robust against weak disorder. Intriguingly, disorder itself can also induce topological phases–exemplified by the Anderson topological insulator in noninteracting systems 5. Experimental studies on disorder induced topology have so far been limited to band topology 6, 7, 8, 9, 10, 11, 12, 13, 14. Here we report direct observations of structural disorder induced many-body interacting topological phase in an atom array at half-filling, whereby random offsets to tweezer locations forming a lattice implement structural disorder, causing fluctuating long-range dipolar interactions between tweezer confined single atoms. The ground state degeneracy in disordered configurations is detected and compared to a regular lattice. The induced topological phase is vindicated by the spatially resolved atom-atom correlation functions for different forms of dimer compositions. By probing the quench dynamics of a highly excited state, we observe markedly slower decay of edge spin magnetization in comparison to the bulk spin one, consistent with the presence of topologically protected edge modes in disordered lattices. Our experiments open a new direction for studying the interplay between structural disorder and strongly interacting topological matter in Rydberg atom arrays.
Topological phases of matter constitute a novel class of quantum states that extends beyond the conventional symmetry-breaking paradigm. Since disorder is ubiquitous in either natural or synthetic systems, understanding its influence on topological phases is of fundamental interest 15, 16. Typically viewed as detrimental to ordered phase, disorder can in fact also induce topological phase–for example, in the celebrated topological Anderson insulator 5. This counterintuitive phenomenon has motivated experimental studies across a range of systems, from the classical ones of waveguide arrays 6, 11, 12, photonic crystals 8, 13, and acoustics 9, 10, to the quantum ones of atomic condensates in momentum lattices 7 and superconducting circuit 14. However, all realizations are concentrated so far on noninteracting particles or fully captured by band topology of single-particle states.
@@
Discussion and outlook.—Our work opens a broad avenue to study the interplay between structural disorder and topology in a fully tunable quantum platform. At the single-particle level, future explorations could address how structural disorder affects higher-dimensional topological phases, such as the Chern insulators 55, 56 and higher-order topological phases 57, 34. At the many-body level with interactions, it is now possible to examine many-body topological phenomena in higher dimensions, including spin liquids in two-dimensional amorphous lattices 58, 59. Furthermore, the Rydberg atom array platform can be employed to explore amorphous Floquet topological phases 60 by incorporating periodic driving. In addition, it also offers new opportunities to study the interplay between localization and topology, such as topologically protected quantum dynamics in disordered systems 52.
@@ Acknowledgements.
We acknowledge significant help and enlightening discussions with Cheng Chen and Songtao Huang, and helpful contributions by Yuanjiang Tang, Chao Liang, Xiangliang Li, Xiaoling Wu, Boyang Wang, and Yifan Wang in the early stages of building up the experimental platform. We acknowledge ChatGPT for language polishing. DMRG calculations were performed using the TeNPy tensor network library 61. We finally acknowledge the support by Center of High Performance Computing, Tsinghua University. This work is supported by the Innovation Program for Quantum Science and Technology (2021ZD0302100). LY is also supported by NSFC (Grants No. 12361131576 and No. 92265205). MKT is supported by NSFC (Grants. No.12234012 and W2431002). YFM, KL, and YX are supported by NSFC (Grants. No.12474265 and 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604).
@@
Previous studies have shown that the inversion symmetry can protect the topology in 1D 62, 63. We now present a proof showing that, under this symmetry, the polarization $P_{\mathrm{S}}$ can only take the values of either 00 or $0.5$. Specifically, consider the Hamiltonian $H^{\textrm{S}}$ that respects the inversion symmetry, i.e., $[H^{\textrm{S}},U_{R}]=0$, where $U_{R}$ is the inversion operator. When acting on $|i\rangle$, the operator yields $U_{R}|i\rangle=|\mathcal{R}(i)\rangle$, where $\mathcal{R}(i)$ denotes the site index of the inversion partner of site $i$. This symmetry ensures that $|u_{i}\rangle$ in $U$ is also an eigenstate of $U_{R}$ corresponding to an eigenvalue $\nu_{i}$ that takes a value of $1$ or $-1$. With this in mind, we can derive that
@@
For bosonic SPT phases, inversion symmetry plays the role of time-reversal symmetry based on the crystalline equivalence principle 64, 65, leading to a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ classification in 1D 66, provided the system also respects the $U(1)$ symmetry. Here we will present a proof demonstrating that inversion symmetry can protect the quantization of $P_{\mathrm{M}}$ in the spin model. Note that an alternative proof has been provided for a spin-1 model in Ref. 50.
@@
In the above proof, we have made use of the fact that $\sigma^{z}$ is conserved corresponding to the $U(1)$ symmetry. As such, the symmetry class of the system is given by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$, where the spatial inversion symmetry is treated as an antiunitary $\mathbb{Z}_{2}$ internal symmetry through the crystalline equivalence principle 64, 65. Based on previous classification results 35, the average SPT phase protected by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$ in 1D is given by

### v2

@@ Abstract
Topological phases of matter can appear in noninteracting systems, as in band topology, or interacting systems, such as in spin models, with their defining features typically robust against weak disorder. Intriguingly, disorder itself can also induce topological phases–exemplified by the Anderson topological insulator in noninteracting systems. Experimental studies on disorder induced topology have so far been limited to band topology. Here we report direct observations of structural disorder induced many-body interacting topological phase in an atom array at half-filling, whereby random offsets to tweezer locations forming a lattice implement structural disorder, causing fluctuating long-range dipolar interactions between tweezer confined single atoms. The ground state degeneracy in disordered configurations is detected and compared to a regular lattice. The induced topological phase is also vindicated by the spatially resolved atom-atom correlation functions for different forms of dimer compositions. By probing the quench dynamics of a highly excited state, we observe markedly slower decay of edge spin magnetization in comparison to the bulk spin one, consistent with the presence of topologically protected edge modes in disordered lattices. Our experiments open a new direction for studying the interplay between structural disorder and strongly interacting topological matter in Rydberg atom arrays.
Topological phases of matter constitute a novel class of quantum states that extends beyond the conventional symmetry-breaking paradigm 1, 2, 3, 4. Since disorder is ubiquitous in either natural or synthetic systems, understanding its influence on topological phases is of fundamental interest 5, 6. Typically viewed as detrimental to ordered phase, disorder can in fact also induce topological phase–for example, in the celebrated topological Anderson insulator 7. This counterintuitive phenomenon has motivated experimental studies across a range of systems, from the classical ones of waveguide arrays 8, 9, 10, photonic crystals 11, 12, and acoustics 13, 14, to the quantum ones of atomic condensates in momentum lattices 15 and superconducting circuit 16. However, all realizations are concentrated so far on noninteracting particles or fully captured by band topology of single-particle states.
@@
Discussion and outlook.—Our work opens a broad avenue to study the interplay between structural disorder and topology in a fully tunable quantum platform. At the single-particle level, future explorations could address how structural disorder affects higher-dimensional topological phases, such as the Chern insulators 55, 56 and higher-order topological phases 57, 34. At the many-body level with interactions, it is now possible to examine many-body topological phenomena in higher dimensions, including spin liquids in two-dimensional amorphous lattices 58, 59. Furthermore, the Rydberg atom array platform can be employed to explore amorphous Floquet topological phases 60 by incorporating periodic driving. In addition, it also offers new opportunities to study the many-body localization induced by structural disorder 61 and the interplay between localization and topology, such as topologically protected quantum dynamics in disordered systems 52.
Note added. During the completion of this work, we became aware of related work studying average symmetry protected topology in an optical lattice system 62.
@@ Acknowledgements.
We acknowledge significant help and enlightening discussions with Cheng Chen and Songtao Huang, and helpful contributions by Yuanjiang Tang, Chao Liang, Xiangliang Li, Xiaoling Wu, Boyang Wang, and Yifan Wang in the early stages of building up the experimental platform. We acknowledge ChatGPT for language polishing. DMRG calculations were performed using the TeNPy tensor network library 63. We finally acknowledge the support by Center of High Performance Computing, Tsinghua University. This work is supported by the Innovation Program for Quantum Science and Technology (2021ZD0302100). LY is also supported by NSFC (Grants No. 12361131576 and No. 92265205). MKT is supported by NSFC (Grants. No.12234012 and W2431002). YFM, KL, and YX are supported by NSFC (Grants. No.12474265 and 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604).
@@
Previous studies have shown that the inversion symmetry can protect the topology in 1D 64, 65. We now present a proof showing that, under this symmetry, the polarization $P_{\mathrm{S}}$ can only take the values of either 00 or $0.5$. Specifically, consider the Hamiltonian $H^{\textrm{S}}$ that respects the inversion symmetry, i.e., $[H^{\textrm{S}},U_{R}]=0$, where $U_{R}$ is the inversion operator. When acting on $|i\rangle$, the operator yields $U_{R}|i\rangle=|\mathcal{R}(i)\rangle$, where $\mathcal{R}(i)$ denotes the site index of the inversion partner of site $i$. This symmetry ensures that $|u_{i}\rangle$ in $U$ is also an eigenstate of $U_{R}$ corresponding to an eigenvalue $\nu_{i}$ that takes a value of $1$ or $-1$. With this in mind, we can derive that
@@
For bosonic SPT phases, inversion symmetry plays the role of time-reversal symmetry based on the crystalline equivalence principle 66, 67, leading to a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ classification in 1D 68, provided the system also respects the $U(1)$ symmetry. Here we will present a proof demonstrating that inversion symmetry can protect the quantization of $P_{\mathrm{M}}$ in the spin model. Note that an alternative proof has been provided for a spin-1 model in Ref. 50.
@@
In the above proof, we have made use of the fact that $\sigma^{z}$ is conserved corresponding to the $U(1)$ symmetry. As such, the symmetry class of the system is given by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$, where the spatial inversion symmetry is treated as an antiunitary $\mathbb{Z}_{2}$ internal symmetry through the crystalline equivalence principle 66, 67. Based on previous classification results 35, the average SPT phase protected by $U(1)\times(\mathbb{Z}_{2}^{T})^{\mathrm{avg}}$ in 1D is given by